# Frontier AI’s Impact on the Cybersecurity Landscape

### v2

@@ 3.1.1 Qualitative impact analysis
Figure 2 shows frontier AI’s qualitative impact on attack steps. Most attack steps show frontier AI’s research influence, with some demonstrating large-scale real-world effects.
Figure 2: A comprehensive categorization of frontier AI’s marginal risks to existing cyberattacks with qualitative impact assessment. Heatmap colors indicate impact levels: Green–no demonstrated effect, Yellow–demonstrated in research papers, Orange–demonstrated in real world, Red–large-scale real world impact. The figure reveals AI’s widespread application in achieving final attack goals but limited use in specific attack stages. This disparity exists because most reports disclose only final impacts rather than attack procedures, obscuring AI’s specific contributions. For precision, we include only reports with clear specifications for individual attack steps.
@@
Attack step 1: Reconnaissance. Recent research explored frontier AI in multiple reconnaissance methods: active scanning [270], victim information gathering [73], and open-source database search [231]. In addition, real-world attacks increasingly use LLMs for active scanning (e.g., satellite capability understanding by Forest Blizzard) and information gathering about victims (e.g., by Salmon Typhoon) [200].
Attack step 2: Weaponization. Recent research demonstrated LLMs can generate functional Windows PE malware with high evasion rates against VirusTotal [213] ($>$70% for Ransomware, Worm, and DDoS). Real-world reports of AI-generated malware primarily focus on simple attacks like DDoS [23, 281]. Although LLMs are aligned to reject harmful queries, these alignment mechanisms can be evaded through jailbreaking [72, 187]. There are also early works on fine-tuning LLMs for attack generation, such as mutating malware to evade detection [248].
@@
Research has leveraged LLMs for vulnerability identification across various network and software systems [253, 186, 174, 240, 80]. Real-world attacks have also begun employing LLMs for vulnerability detection, including MSDT vulnerability (CVE-2022-30190) [200] and SQLite overflow bugs [119]. Moreover, research showed LLMs can generate PoCs for memory corruption bugs in C/C++ [286, 311]. AI agent teams can exploit real-world N-day [90] and zero-day [91, 215] vulnerabilities (e.g., SQL/XSS injection, concurrency attacks, remote code execution) and solve CTF challenges [271, 329, 251]. However, the study on exploitation remains mainly at a research level, with only a few public real-world reports [142, 200, 295].
@@
Attack step 3-5: Delivery & exploitation & installation. Initial access and installation have been demonstrated in the real world [208, 142] while persistence remains at a research level [245]. Initial access is achieved through compromised accounts or infrastructure [96, 208, 142], with AI assisting in password guessing [135, 288, 132] and bypassing defenses, including intrusion detection, CAPTCHA, and two-factor authentication [222, 74, 200, 70]. LLMs also assist malware configuration and installation [105]. Research works also used AI to maintain persistence across system restarts and credential changes, e.g., LLMs help masquerading by disguising malicious activities as benign [245, 28].
@@
Attack step 6: Commend and control (C&C). Research demonstrates frontier AI can facilitate privilege escalation [121, 68, 158, 122]. For instance, ChainReactor [68] used LLM planning to generate exploit chains for privilege escalation. For stable and stealthy C&C channels, attackers can employ LLMs for automated domain generation, creating resilient remote management of compromised systems [245, 13, 28]. Recent research showed LLMs can assist with adaptation [57], allowing self-learning malware to create environment-tailored attacks. Despite these research advances, only a few real-world reports exist for credential access in this step [148, 244, 211].
@@
Attack step 5: Action on objectives. As shown in Figure 2, AI-enhanced cyberattacks have proliferated across web, IoT, mobile, and cloud platforms, causing real-world financial losses, privacy leakage, data manipulation, and resource hijacking [23, 137, 281, 142, 96, 208, 63, 327, 99, 36, 316, 260]. A recent attack deployed AsyncRAT malware using LLMs to leak sensitive information [137]. Frontier AI has amplified attacks against web applications and IoT devices in eCommerce, including business logic abuse, API violations, bad bot attacks, and targeted DDoS on retail sites [191, 100]. These attacks have disrupted operations, stolen data, and exploited digital engagement for financial gain [142]. CrowdStrike reports a 110% rise in cloud-aware cyberattacks and 75% increase in intrusions year-over-year with AI enhancement [63]. For instance, threat actor INDRIK SPIDER exploited cloud-based credential storage, accessed Azure Key Vault credentials, and even used ChatGPT while managing Azure resources [63].
There are some early research exploring end-to-end attacks using LLMs or AI agents [306, 289, 280, 260]. SecCodePLT [316] used LLMs to generate end-to-end attacks with pre-specified attack steps and prompts. Autoattacker [306] proposed a multi-agent system with planning and generation agents to automatically analyze attack goals and generate corresponding attacks. They demonstrated that AI agents can plan and generate attacks for well-defined attack goals in a controlled environment.
Future impact. First, frontier AI could potentially lower the attack thresholds, particularly for attack steps 1&2, where it assists with information gathering even with limited access. Frontier AI models are trained on vast datasets, which potentially allows them to guess and reason partial information. For instance, deep learning models can achieve high accuracy in reverse-engineering stripped binaries [114, 220, 305]. In addition, as foundation models advance, they are expected to better understand program logic and semantics, likely enabling them to identify and exploit logical bugs under partial information (demonstrated by the emerging reasoning models [112, 203, 169]).
Second, we envision frontier AI can make attacks more automated and stealthy. For example, automated exploit chain generation in the attack step 3 (exploitation and installation) has been challenging. Previous approaches like counter-example-guided program synthesis [92] and rule-based methods [301] require expertise and significant manual effort for complex programs. Frontier AI offers solutions to overcome these limitations and other labor-intensive attacks, e.g., resource hijacking and operational disruption. Third, AI’s continuous learning capabilities enable automatic evolution and customized attack generation. By learning from both successful and failed attempts, attackers can refine strategies for more efficient and stealthy attacks.
[…]
                                         Attack step Coverage Metric Dynamic updates
[…]
3.1.2 Quantitative impact analysis
Table II summarizes existing benchmarks for evaluating LLMs and AI agents in different attack steps. In summary, except for vulnerability exploitation in step 2, existing benchmarks are limited in risk coverage, metric design, and dynamic evolution. Table VI in the Appendix details SOTA LLMs’ performances on existing attack benchmarks.
Attack step 1: Reconnaissance. Few benchmarks exist for reconnaissance. Most of them [274, 184, 286] evaluate LLMs through text-based reconnaissance questions and use another LLM as the judge. They fail to test the capabilities of generating and executing actual attacks. AutoPenBench [101] included reconnaissance but covers limited methods and scenarios. Results show that simple AI agents (LLMs with tool uses) demonstrate early success in reconnaissance tasks like discovering potential victims on remote networks or conducting surveillance. However, agents cannot identify target-specific vulnerable services due to limited service information from standard tools (e.g., Nmap) and the lack of deeper interactions with target systems.
Attack step 2: Weaponization. For malware creation, RedCode [111] benchmarked a prompt set that queries LLMs to generate malware in Python and uses VirusTotal as the judge. Another work crafted a benchmark to test the safety alignment of LLMs with jailbreaking prompts on attack generation [285]. They showed that Claude 3.5 Sonnet is more resilient than GPT-4o and Genimi-1.5. Regarding vulnerability exploitation, recent works designed some benchmarks for crash and PoC generation (e.g., CyberSecEval [286] and SecCodePLT [316]). They crafted specific coding tasks that may encounter known vulnerabilities in the CWE database and evaluate where LLMs generate vulnerable code when finishing these tasks. SecCodePLT [316] supports dynamic testing, which is more precise than LLM judgment. Their results showed that both SOTA commercial and open-source models frequently generate vulnerable code, with stronger models being more vulnerable due to their enhanced task understanding and execution.
While no dedicated exploit generation benchmark exists, existing CTF benchmarks [251, 329] and model pre-deployment tests [279] included sub-benchmarks evaluating exploitation capabilities. Their findings indicate SOTA commercial models like OpenAI-o1 and Claude 3.5 Sonnet achieve skill levels matching technical non-experts and cybersecurity apprentices in vulnerability exploitation. Claude 3.5 Sonnet can exploit more complex vulnerabilities at cybersecurity practitioner level. However, SOTA models rarely solve expert-level challenges, e.g., OpenAI-o1 has a very low success rate [329].
Attack step 3$\sim$7. Benchmarks for these later attack steps remain severely limited in coverage and quality. Similar to Reconnaissance, most existing benchmarks employ text-based queries rather than generating actual attacks [274, 184, 286, 156]. SecCodePLT [316] and AutoPenBench [101] include partial attack steps and use dynamic testing as the judge. SecCodePLT [316] provides a network system with pre-specified attack paths and targets, enabling end-to-end attacks, including reconnaissance, initial control, C2, and collection. Their results showed that Claude 3.5 Sonnet’s safety alignment rejects most attack generation queries, while GPT-4o generates end-to-end attacks with low success rates. CTF benchmarks [251, 329] include pwn challenges testing privilege escalation, showing that commercial LLMs slightly outperform open-source ones.
Future efforts. Existing benchmarks are limited across multiple dimensions. Constructing practical, comprehensive, high-quality benchmarks that can be dynamically updated requires significant community-wide effort. Section 7.1 provides actionable recommendations for achieving this goal.
Enhancing human attackers. Frontier AI not only enables fully automated attacks but also enhances human attackers’ productivity. At a high level, attackers can leverage AI to automate partial attack steps or provide initial attack solutions. For example, attackers can leverage AI to generate exploits for individual vulnerabilities and then chain them together manually. They can also further polish the LLm-generated phishing attacks to be more stealthy and convincing. While frontier AI will improve attack automation, we anticipate that in the short term, there will be more real-world attacks leveraging such human-AI collaborations.
Figure 3: Marginal risks of frontier AI to existing attacks against humans.
@@ 3.2 Attacks Targeting Humans
Humans remain the weakest link in security systems, with AI-driven attacks against them proliferating. Figure 3 shows that most human-targeted attacks have demonstrated large-scale impacts in the real world. Studies reveal a 135% increase in social engineering attacks following ChatGPT’s widespread adoption in 2023 [67]. Phishing attacks constitute a significant portion [43, 89, 30], with voice phishing surging 260% in 2023 compared to 2022 [262]. While studies from 2023 and mid-2024 [43, 88] found AI-generated phishing less effective than human-crafted attacks, recent research [128, 17] demonstrates fully AI-automated phishing can now match human expert effectiveness.
@@
Deepfakes represent one of the most serious threats. Hong Kong police documented attackers using AI-driven deepfake technology to impersonate a chief financial officer during a video call, successfully defrauding \$25 million [49]. Child exploitation constitutes another severe AI misuse, including AI-generated child sexual abuse material, automated grooming, and detection evasion techniques [266]. AI-driven bots enable cyberbullying and harassment [134], facilitating emotional manipulation and behavioral control. Some nations reportedly employ AI for cyber espionage and psychological manipulation [238]. Misinformation presents another critical concern, with research demonstrating AI’s growing capabilities to influence human opinions and decisions [223, 21, 16, 104, 62]. These highlight how these capabilities can be exploited to spread misinformation and manipulate human behavior at scale.
Future impact. The impact of frontier AI attacks against humans will intensify dramatically. Advancing AI systems will enable more convincing, stealthy, and targeted social engineering attacks that bypass traditional security measures and human judgment. Multi-modal AI capabilities—combining text, image, audio, and video—will blur distinctions between authentic and fraudulent content, resulting in greater social damage from deepfake fraud and more effective public opinion manipulation through realistic synthetic media.
Key takeaways. Qualitative: Real-world AI-facilitated attacks have been reported across various systems, yet AI involvement in specific attack steps remains at an early stage and is limited to research settings. Steps 1&2 (reconnaissance and weaponization) show greater AI integration compared to later steps, with notable advances in malware creation and vulnerability exploitation.
Quantitative: Benchmarks show AI cannot generate valid malware but can exploit bugs at cybersecurity apprentice level. GPT-4o can generate end-to-end attacks, while Claude 3.5 Sonnet rejects most attack-related queries. Attack benchmarks can be improved in attack step coverage, metrics, and dynamic updates. Prediction: Frontier AI will lower the thresholds for attack steps 1 and 2 by reasoning with partial information, as well as enabling more automated and stealthy end-to-end attacks. Human: Frontier AI attacks already have a large-scale impact through social engineering, fraud, deepfakes, child abuse, and misinformation.
4 Frontier AI’s Impact on Defenses
This section discusses frontier AI’s impact on defenses for systems and humans, where we break down system defenses into empirical defenses and those with provable guarantees.
@@ 4.1.1 Qualitative analysis
Figure 4 shows our categorization of cyber defenses summarized from existing literature [261, 93, 124]. As shown in the figure, frontier AI has been predominantly utilized in proactive testing and attack detection but lacks widespread adoption in vulnerability triage and remediation. Currently, vulnerability remediation remains a labor-intensive and time-consuming process.
Defense step 1: Proactive testing include penetration testing and proactive vulnerability detections. Early research works automate penetration testing with LLMs and AI agents [155, 73, 120, 306, 273, 60, 133, 145] and more recent ones focused on optimizing agents’ workflows and tool calls [155, 73]. Due to the dual usage nature, these methods may also be used to facilitate attacks. LLMs have been used in industries for penetration testing [106] (But we lack statistics on whether they have been widely used).
@@
More and more research leverages AI to facilitate vulnerability detection. This first includes developing customized graph neural networks [338, 307] and uses general LLMs and fine-tuned ones to directly identify vulnerabilities in various software applications [80, 5, 204, 240, 186, 174, 254, 224, 84, 269]. These works reported mixed results when comparing AI with traditional static analysis tools (e.g., CodeQL, Snyk, and Fortify) and are all concerned about the errors of AI and limited scalability to repository-level detection involving multiple functions and files. Early evidence showed that AI agents can help with vulnerability detection (e.g., Google Project Zero’s report on a real-world vulnerability: stack buffer underflow in SQLite) [327, 170]. Some research works also improved dynamic fuzzing with frontier AI [139, 253, 310, 207, 196, 76, 330], such as LLM-based pattern/rule/seed generation and RL-based power scheduling [255]. There are also early explorations in using LLMs for hybrid methods that combine static with dynamic program analysis, i.e., HyLLfuzz [195] developed an LLM-based concolic execution that uses LLMs as the solver for finding inputs to specific program branches.
Future impact. First, frontier AI could enhance program-analysis-based vulnerability detection approaches, better balancing soundness, completeness, and scalability. Specifically, AI agents’ scheduling and planning capabilities can improve dynamic fuzzing’s search efficiency and help select promising states for symbolic execution to prevent path explosion. These advancements can be extended to concolic execution and other hybrid approaches. However, frontier AI’s lack of robustness may offset marginal benefits, necessitating in-depth explorations that can design AI-based solutions for tackling traditional methods’ key challenges (see Section 7). Second, reducing AI-based vulnerability detection error rates requires improving models’ vulnerability reasoning capabilities. This requires high-quality datasets with additional reasoning data, such as vulnerable code with intermediate program states and human reasoning processes. Such supplementary information is crucial because frontier LLMs can process intermediate data to enhance decision-making beyond simple classification. Then, we need to design customized training approaches that improve the models’ reasoning capabilities, which help models handle ambiguous and complex cases (evidenced by recently emerged reasoning models). Third, with stronger models as agents, frontier AI could handle more complex cases for real-world code base.
Figure 4: Impact of frontier AI on different defense steps. It also shows the relations with the attack steps in Figure 2.
@@
Defense step 2: Attack detection. Malware detection and network intrusion detection are two applications where ML has been widely used [290, 302, 45, 283, 161]. Recent research showed that frontier AI can address key challenges of traditional AI-based attack detections: reliance on feature engineering and labeled data, and poor generalizability. First, traditional AI models like decision trees and linear regression rely heavily on well-designed feature engineering, while deep learning can learn directly from raw inputs. When being used for attack detection, deep learning also eliminates the need for feature engineering and can explore large learning spaces beyond human-designed features, potentially leading to more effective solutions.
@@
Second, existing AI-based attack detection requires large amounts of well-labeled data, which is challenging for security data [12, 302]. Transformers can learn general knowledge from vast unsupervised data [282, 77, 35], which can be customized for specific downstream tasks during supervised fine-tuning, significantly reducing reliance on large labeled datasets [82, 77]. Recent works showed that transformers trained on raw network traffic bytes outperform supervised models trained on handcrafted features in intrusion [117, 178] and malware detection [7, 324]. Third, existing AI-driven attack detection faces generalizability issues with out-of-distribution (OOD) data. Frontier AI models, trained on billions of data points or trillions of tokens, encompass wider distribution ranges during training. This reduces OOD likelihood during inference, which improves the generalizability in intrusion detection [117, 320, 98, 265] and malware detection [7, 324].
Future impact. Frontier AI-driven attack detections can continue improving in generalizability, particularly to attack evolvement and unseen threats, as well as adversarial robustness. As the models increase in capacity and reasoning capability, they could be better at finding hidden connections between existing attacks and evolved ones. In addition, through lifelong learning, these models can continuously update themselves to stay current with new threats. Furthermore, with enhanced safety alignment and more powerful and specialized guardrails, frontier AI models can become increasingly robust in threat detection. Recent advancements in model interpretability [340] allow developers to access a model’s internal representations and reasoning processes when a decision is made, which improves model interpretability and helps identify attacks. Finally, frontier AI could also be used for attack monitoring and instant alerting.
@@
Defense step 3: Triage & forensic. First, for PoC generation, existing works explored improving dynamic fuzzing and symbolic execution with LLMs [292, 126]. For example, recent research [292] proposed using LLMs to translate Python path constraints into Z3 code, which enabled symbolic execution on Leetcode problems with complex control flows and list data structures. Second, for root cause analysis, traditional methods [114, 202, 64] employ static and dynamic analysis tools, such as reverse execution, backward taint analysis, value-set analysis, and alias analysis. Some recent works explored developing AI agents for root cause analysis [235, 308, 225]. These agents collect additional information (e.g., logs) through tool calling and leverage advanced prompting techniques (e.g., ReAct [317] and SelfReflection) to improve root analysis performance. Later in Section 4.1.3, we discuss how frontier AI facilitates SOC, which is also related to triage and forensics. Third, for reverse engineering, early works used recurrent networks for function boundary and signature recovery [257, 56]. However, such models have limited generalizability. Recent research [220, 305] explored pre-training or fine-tuning foundation models for binary codes and demonstrated better performance than RNN-based models and traditional rule-based tools in multiple reverse engineering tasks.
Future impact. Frontier AI can address key challenges in existing heuristic-based approaches for triage and forensic. For example, alias analysis is crucial for recovering data flow, especially with partial execution traces. Research demonstrates recurrent networks’ potential for alias analysis [114], and transformer-based models could further enhance this capability. Frontier AI can also enhance PoC generation and reverse execution as foundation models advance in capability, generalizability, and robustness. More broadly, we believe AI agents will become a promising approach that combines AI-based triage with traditional methods. Specifically, we can build multi-agent systems for root cause analysis, integrating existing techniques like reverse execution, taint analysis, and value-set analysis with AI-driven alias analysis as tools. In such a system, LLMs would drive the analysis by generating step-by-step plans, calling tools, and dynamically adjusting the analysis logic.

### v3

@@ 4.1 Attacks Targeting Systems
AI’s influence on most steps remains at research, though several demonstrate large-scale real-world effects (Figure 2).
@@
Attack Step 1: Reconnaissance. Real-world attacks increasingly use LLMs and AI agents for active scanning (e.g., satellite capability understanding by Forest Blizzard) and information gathering about victims (e.g., by Salmon Typhoon) [160]. Recent research explores frontier AI in multiple reconnaissance methods: active scanning [221], victim information gathering [56], and open-source database search [189].
@@
Attack Step 2: Weaponization. Malware generation mainly stays at the research level, where real-world incidents mainly report DDoS attacks [20, 230]. Specifically, research shows that LLMs can generate Windows PE malware with high evasion rates against VirusTotal [173] (>>70% for Ransomware, Worm, and DDoS); LLM alignments can be evaded [55, 150]; LLMs are also fine-tuned to generate attacks [201]. For vulnerability identification, LLMs are used in various network and software systems [205, 149, 135, 194, 63, 143]. Real-world attacks also begin employing LLMs for vulnerability detection, including MSDT vulnerability (CVE-2022-30190) [160] and SQLite overflow bugs [93]. Moreover, research shows LLMs can generate PoCs for memory corruption bugs in C/C++ [236, 255, 242]. AI agents can exploit real-world N-day [71, 270] and zero-day [72, 175, 282] vulnerabilities (e.g., SQL/XSS injection, concurrency attacks, remote code execution) and solve CTF challenges [223, 272, 203]. However, the study on exploitation remains mainly at a research level, with only a few public real-world reports [115, 160, 245].
@@
Attack Steps 3-5: Delivery & exploitation & installation. Initial access and installation have been demonstrated in the real world [169, 115], while persistence remains at a research level [199]. Initial access is achieved through compromised accounts or infrastructure [75, 169, 115], with AI assisting in password guessing [107, 238, 104] and bypassing defenses, including intrusion detection, CAPTCHA, and two-factor authentication [181, 57, 160, 53]. LLMs and AI agents can also assist malware configuration and installation [83]. Research works employ frontier AI to maintain persistence across system restarts and credential changes, e.g., AI helps masquerading by disguising malicious activities as benign [199, 23].
@@
Attack Step 6: Commend and control (C&C). Research has demonstrated frontier AI can facilitate privilege escalation [95, 52, 125, 96, 227]. For instance, ChainReactor [52] uses LLM planning to generate exploit chains for privilege escalation. For stable and stealthy C&C channels, attackers can employ LLMs for automated domain generation, creating resilient remote management of compromised systems [199, 11, 23]. Recent research shows LLMs can assist with adaptation [43], allowing self-learning malware to create environment-tailored attacks. Only a few real-world reports exist for credential access in this step [119, 198, 171].
@@
Attack Step 7: Action on objectives. AI-enhanced cyberattacks have proliferated across web, IoT, mobile, and cloud platforms, causing real-world financial losses, privacy leakage, data manipulation, and resource hijacking [20, 109, 230, 115, 75, 169, 48, 269, 78, 30, 260, 211]. A recent attack deploys AsyncRAT malware using LLMs to leak sensitive information [109]. Frontier AI amplifies attacks against web applications and IoT devices in e-commerce, including business logic abuse, API violations, bad bot attacks, and targeted DDoS on retail sites [152, 79]. These attacks disrupt operations, steal data, and exploit digital engagement for financial gain [115]. CrowdStrike reports a 110% rise in cloud-aware cyberattacks and 75% increase in intrusions year-over-year with AI enhancement [48]. Early research explores end-to-end attacks using LLMs or AI agents [250, 239, 229, 211].
@@ 4.2 Attacks Targeting Humans
We find that most human-targeted attacks have demonstrated significant real-world impacts (Figure 2). Studies show a 135% increase in advanced social engineering attacks following ChatGPT’s release from Jan to Feb 2023 [51]. Phishing attacks form a significant portion [32, 70, 25], with voice phishing surging 260% in Q4 2023 compared to Q4 2022 [213]. While studies from 2023 and mid-2024 [32, 68] find AI-generated phishing less effective than human-crafted attacks, recent research [101, 14] demonstrates AI-automated phishing can now match human expert effectiveness.
@@
Deepfakes represent one of the most serious threats. Hong Kong police documented attackers using AI-driven deepfake technology to impersonate a chief financial officer during a video call, successfully defrauding \$25 million [36]. Child exploitation constitutes another severe AI misuse, including AI-generated child sexual abuse material, automated grooming, and detection evasion techniques [218]. AI-driven bots enable cyberbullying and harassment [106], facilitating emotional manipulation and behavioral control. Some nations reportedly employ AI for cyber espionage and psychological manipulation [192]. Misinformation presents another critical concern, with research demonstrating AI’s growing capabilities to influence human opinions and decisions [182, 18, 13, 82, 47].
@@ 4.3 Proactive and Reactive Defenses
We find that frontier AI has been predominantly utilized in proactive testing and attack detection but lacks widespread adoption in vulnerability triage and remediation (Figure 2). Especially, vulnerability remediation remains a labor-intensive and time-consuming process.
@@
Defense Step 1: Proactive testing include penetration testing and proactive vulnerability detections. Early research works automate penetration testing with LLMs and AI agents [123, 56, 94, 250, 224, 45, 105, 116] and more recent ones focus on optimizing agents’ workflows and tool calls [123, 56]. More and more research leverages AI to facilitate vulnerability detection. This first includes developing customized graph neural networks [281, 252] and uses general LLMs and fine-tuned ones to directly identify vulnerabilities in various software applications [63, 4, 163, 194, 149, 135, 206, 183, 65, 220]. These works report mixed results when comparing AI with traditional static analysis tools (e.g., CodeQL, Snyk, and Fortify) and are all concerned about the errors of AI and limited scalability to repository-level detection involving multiple functions and files. To mitigate these weaknesses and capitalize on LLMs’ strong reasoning capabilities, recent work proposes hybrid workflows that integrate LLMs with static analysis tools for repository-level vulnerability detection [130, 136]. Early evidence shows that AI agents can help with vulnerability detection (e.g., Google Project Zero’s report on a real-world vulnerability: stack buffer underflow in SQLite) [269, 89]. Some research works also improve dynamic fuzzing with frontier AI [112, 205, 254, 168, 156, 58, 273, 275], such as LLM-based pattern/rule/seed generation and RL-based power scheduling [207]. There are also early explorations in using LLMs for hybrid methods that combine static with dynamic program analysis, i.e., HyLLfuzz [155] develops an LLM-based concolic execution that uses LLMs as the solver for finding inputs to specific program branches.
@@
Defense Step 2: Attack detection. Malware detection and network intrusion detection are two applications where ML has been widely used [240, 247, 34, 233, 127, 144]. Recent research shows that frontier AI can address key challenges of traditional AI-based attack detection: reliance on feature engineering and labeled data, and poor generalizability. First, traditional AI models rely heavily on well-designed feature engineering, while deep learning can learn directly from raw inputs. When being used for attack detection, deep learning also does not need feature engineering and can explore large learning spaces beyond human-designed features, potentially leading to more effective solutions.
@@
Second, existing AI-based attack detection requires large amounts of well-labeled data, which is challenging for security data [10, 247]. Transformers can learn general knowledge from vast unsupervised data [231, 59, 29], significantly reducing reliance on large labeled datasets [64, 59]. Recent works show that transformers trained on raw network traffic bytes outperform supervised models trained on handcrafted features in intrusion [91, 140] and malware detection [6, 266].
Third, existing AI-driven attack detection faces generalizability issues with out-of-distribution (OOD) data. Frontier AI models, trained on billions of data points or trillions of tokens, encompass wider distribution ranges during training. This reduces OOD likelihood during inference, which improves the generalizability in intrusion detection [91, 262, 77, 216] and malware detection [6, 266].
@@
Defense Step 3: Triage & forensic. First, for PoC generation, existing works explore improving dynamic fuzzing and symbolic execution with LLMs [241, 99], demonstrating the strong potential of AI agents [242, 166]. For example, recent research [241] proposes using LLMs to translate Python path constraints into Z3 code, which enabled symbolic execution on Leetcode problems with complex control flows and list data structures. FaultiLine [166] introduces an agent that performs dataflow tracing, branch condition reasoning, and test case refinement to generate and validate PoCs for target vulnerabilities. Second, for root cause analysis, traditional methods [90, 162, 50] employ static and dynamic analysis tools, such as reverse execution, backward taint analysis, value-set analysis, and alias analysis. Some recent works explore developing AI agents for root cause analysis [191, 253, 184]. These agents collect additional information (e.g., logs) through tool calling and leverage advanced prompting techniques to improve root analysis performance. Third, for reverse engineering, early works use recurrent networks for function boundary and signature recovery [209, 42]. However, such models have limited generalizability. Recent research [179, 248, 117, 251, 110] explores pre-training or fine-tuning foundation models for binary codes and demonstrates better performance than RNN-based models and traditional rule-based tools in multiple reverse engineering tasks.